# Yuehao Bai

### v2

@@ 5 Simulations
where $((X_{i},\epsilon_{i}):1\leq i\leq n)$ are i.i.d., $X_{i}$ and $\epsilon_{i}$ are independent with $X_{i}\sim U[0,1]$, $\epsilon_{id}\sim N(0,1)$, the parameters $\mu_{d}$ are given by $\mu_{1}=0.25$, $\mu_{0}=0$, and $\mu_{d}(\cdot)$ are specified as
@@ Appendix A Proofs of Main Results
$\frac{1}{m}\sum_{1\leq j\leq m}\frac{1}{\binom{k}{2}}\sum_{i<i^{\prime}\in\lambda_{j}}R_{i}R_{i^{\prime}}=\frac{1}{\binom{k}{2}}\sum_{1\leq\gamma_{1}<\gamma_{2}<k}\frac{1}{m}\sum_{1\leq j\leq m}R_{i_{j,\gamma_{1}}}R_{i_{j,\gamma_{2}}}~{}.$
@@
=It follows from Assumption 4.1(b) applied across strata and similar arguments to those used to prove (23)–(24) that

### v3

@@ 5 Simulations
where $((X_{i},\epsilon_{0i},\epsilon_{1i}):1\leq i\leq n)$ are i.i.d., $(X_{i},\epsilon_{0i},\epsilon_{1i})$ are independent with $X_{i}\sim U[0,1]$, $\epsilon_{di}\sim N(0,1)$, the parameters $\mu_{d}$ are given by $\mu_{1}=0.25$, $\mu_{0}=0$, and $\mu_{d}(\cdot)$ are specified as
@@ Appendix A Proofs of Main Results
$\frac{1}{m}\sum_{1\leq j\leq m}\frac{1}{\binom{k}{2}}\sum_{i<i^{\prime}\in\lambda_{j}}R_{i}R_{i^{\prime}}=\frac{1}{\binom{k}{2}}\sum_{1\leq\gamma_{1}<\gamma_{2}\leq k}\frac{1}{m}\sum_{1\leq j\leq m}R_{i_{j,\gamma_{1}}}R_{i_{j,\gamma_{2}}}~{}.$
@@
It follows from Assumption 4.1(b) applied across strata and similar arguments to those used to prove (24)–(25) that